# Natural Language Processing for Dialects of a Language: A Survey

### v2

@@ 5.1.1. Shared tasks
Shared tasks have accelerated past work in dialect identification. These have been primarily led by Workshop on NLP for Similar Languages, Varieties and Dialects, also known as VarDial. (Zampieri et al., 2015) is a 2015 shared task to discriminate between similar languages, including American and British English; and Argentinian Spanish and Castilian Spanish. (Zampieri et al., 2014) is the 2014 version of the task which includes Brazilian Portuguese and European Portuguese in addition to the previous. In 2016 version of the same shared task, (Malmasi et al., 2016) introduced French from Canada and from France to the set. In addition, they also added a subtask for dialect classification of Arabic, namely for the dialects: Modern Standard Arabic (MSA) and the Egyptian (EGY), Gulf (GLF), Levantine (LAV) and North African (NOR) dialects. (Goutte et al., 2016) summarises the findings from the past versions of the shared task from 2014-2016. (Zampieri et al., 2019) present a report of a dialect identification shared task from 2019. Their datasets include dialects of German, Chinese, Romanian along with Cuneiform language identification.
@@ 5.1.2. Datasets
Table  3 shows datasets labeled with different dialects. (Aji et al., 2022) report a dataset for several languages and dialects spoken in Indonesia. They observe that language identification works well for certain dialects (Ngoko-Central dialect of Javanese, for example). The paper also discusses code-mixing and orthography variations in these languages. (Dunn, 2019) reports dialect identification on 14 national varieties of English. They show that cross-domain classification (CommonCrawl versus Twitter) also performs poorly. (Cotterell and Callison-Burch, 2014) present a multi-dialect, multi-genre corpus of news comments and tweets written in dialects of Arabic. The tweets are manually annotated for dialect identification on MTurk. (Ramponi and Casula, 2023a) present a benchmark dataset for dialects of Italian. The benchmark is named as DIATOPIT. There has been recent work on creating a corpus of Norwegian dialect data by (Barnes et al., 2021)(Alshutayri and Atwell, 2018) present a large (200K+ instances) corpus for Arabic dialects and Standard Arabic. The data is sourced largely from tweets but also includes comments from newspapers, and Facebook. The data is also being annotated for dialect identification and contains 24K annotated documents. (Le and Luu, 2023) present a parallel corpus for dialects of Vietnamese.
@@ 5.1.3. Feature-based approaches
We now highlight different kinds of features used for dialect identification.
@@
    Acoustic features(Chitturi and Hansen, 2008) is an early work in dialect identification from audio recordings of podcasts in Australian, UK and American English. They use n-gram-based classifiers on the transcripts of the recordings along with Gaussian mixture model-based classifiers on acoustic features. In the language of today, this would be called a multimodal classifier.
@@
    Phonological features: While acoustic features are derived from audio data, phonological features are based on markers in the written scripts. (Darwish et al., 2014) use lexical, morphological and phonological features in a random forest classifier to detect dialects of Arabic spoken in a geographical region. They also use a lexicon of dialectic Egyptian words.
  3. (3)
    Linguistic features(Doğruöz and Nakov, 2014) present a method to predict dialects of Turkish by using light verb constructions. They use a statistical classifier based on verb-based features (base word, verb order, affixes, etc.) for the task.
The combinations of the above set of features have also been reported. While (Hanani and Naser, 2020) work on the detection of dialects from speech, they also use word-level n-gram features. (Salameh et al., 2018) perform fine-grained dialect identification for 25 dialects of Arabic, using Naïve Bayes classifier and word and character n-grams as features.
Limited pre-deep learning work exists in the context of English.  (Lui and Cook, 2013) is an early work in the detection of dialects of English. Specifically, the paper focuses on Australian, British and Canadian English. Their baseline is the LangID classifier (Lui and Baldwin, 2012) where dialects are treated as individual languages. They experiment with classifiers using features such as n-grams and POS-n-grams. This includes a distribution over function words and those in a vocabulary, akin to a clustering algorithm. (Simaki et al., 2017) also presents statistical models (based on SVM, decision tree and so on) for dialects of English: US, UK, AUS, CAN and so on. They use linguistic, POS-tag-based and lexicon-based features.
@@ 5.1.4. Deep learning-based Approaches
Deep learning-based approaches for dialect classification span three alternatives: train embeddings to reflect dialectic variations, use end-to-end LLMs, or predict dialect as a result of inference over dialect features.
Embeddings in focus(Abdul-Mageed et al., 2018) label tweets with 10 dialects of Arabic. The city is considered the dialectic granularity. The analysis compares dialectic variants by looking at word embeddings of words across different dialects. They use word2vec representations to show how dialectic words are captured. (Goswami et al., 2020) build character-to-sentence embeddings to represent words of different dialects. Unsupervised loss is computed in order to generate clusters of representations. While they also test on language identification, the dialect identification part is done on Swiss German dialect. (Jurgens et al., 2017) use a character-based seq2seq model to map dialects. The models used for language identification are RNNs with GRU. (Criscuolo and Aluisio, 2017) use character n-grams to identify language groups. This is followed by convolutional neural network-based dialect classifiers for each language group.
Fine-tuning LLMs(Ramponi and Casula, 2023a) experiment with multiple models including statistical and neural. The fine-tuned AlBERTo model performs the best among umBERTo, mBERT and XLM-R. (Obeid et al., 2020) present CAMeL: a python toolkit for Arabic language processing. It contains a dialect identifier that gives a distribution over multiple dialects. They use dialectic guidelines provided in  (Elfardy and Diab, 2012).
Detecting dialect features: (Demszky et al., 2020) introduce an approach for dialect classification using a novel multi-task approach that employs dialect feature detection. They train models using two multi-task learning-based approaches using a small number of minimal pairs. They evaluate the output based on 22 dialectical features based on Indian English and demonstrate that such models show the capability of learning to identify features with high accuracy. They show the efficacy of this task by applying it for dialect identification, and by providing a measure of dialect density.
@@ 5.2. Sentiment Analysis
Sentiment analysis is the NLU task of prediction of sentiment polarity of a text. Sentiment analysis encompasses several related tasks such as sarcasm classification and target-specific sentiment analysis. We discuss past work in sentiment analysis along four directions: experiences from annotation (which highlights the challenge of dialects for sentiment analysis), dialect-invariant models, dialect-aware models and finally de-biasing of sentiment analysis models as a post-processing step. Table 4 summarises approaches for sentiment analysis.
@@ 5.2.1. Datasets & Annotation
Several datasets in Arabic sentiment analysis for dialects have been reported such as Moroccan (Oussous et al., 2020) and Levantine (Baly et al., 2019). Dialects can have an impact on annotation itself. (Farha and Magdy, 2022) show that dialect familiarity helps sarcasm annotation. (Mdhaffar et al., 2017) create a dataset of 17000 Facebook comments labeled with sentiment in Tunisian dialect of Arabic. (Assiri et al., 2018) present a sentiment-labeled lexicon of words in the Saudi dialect of Arabic, and use simple counting-based sentiment analysis. (Husain et al., 2022) use weakly supervised labels for sentiment analysis of tweets in the Kuwaiti dialect of Arabic. The labels are then manually validated and updated.
@@ 5.2.2. Dialect-aware representations
(Farha and Magdy, 2022) train BERT-based models for sarcasm detection on data annotated by either of the two groups: those familiar with the dialect and those not. They show that familiarity of dialect improves the quality of the models trained on such a dataset. As a result, representations that capture dialects have been used for sentiment analysis. (Mdhaffar et al., 2017) present models based on SVM and multi-layer perceptron (MLP). (Mulki et al., 2019) use a syntax-ignorant n-gram composition to create embeddings. The classifier model is a dense neural network that works on the addition of word embeddings, with a softmax at the end. (Guellil et al., 2021) propose ‘one’ model for sentiment classification in different dialects of Arabic. They use transliteration to map dialects to Standard Arabic. The sentiment analysis model itself uses word2vec features with statistical classifier. Finally, (Husain et al., 2022) present statistical models based on SVM along with Transformers-based models like BERT.
@@ 5.2.3. Incorporating dialect information in sentiment prediction
(El Mekki et al., 2021) use domain adaptation for sentiment analysis of dialects. Using representations from a BERT encoder, they use two classifiers: sentiment classifier and dialect classifier. The output of the two is later combined for the overall prediction. While this is a two-channel approach, the representation used for the task has also been used to predict dialect of the language. One such example is (Okpala et al., 2022) who present an approach for hate speech detection using African-American English (AAE). In order to do so, they re-train BERT with AAE tweets. Finally adversarial training to regulate debiasing of the hate speech classifier. Specifically, the adversary takes the final representation learned by the hate-speech classifier; and learns to predict the dialect from it. (Kaseb and Farouk, 2022) present a dialect-aware approach for sarcasm detection called the SAIDS model. SAID uses MARBERT to detect dialect and sarcasm. Following that, MARBERT along with sarcasm and dialect output are used to detect sentiment. Evaluated on dialects of Arabic, SAID uses backpropagation only for prediction with respect to the BERT base model. It does not flow through sentiment¡-¿sarcasm or sentiment¡-¿dialect.
@@ 5.2.4. De-biasing sentiment analysis models
Making sentiment analysis agnostic to dialects involves removing dialectic biases in the resultant models. A work of this nature is by (Ball-Burack et al., 2021) who apply adversarial debiasing on resampled data for harmful tweet detection of tweets written in African American English. Resampling of the data uses a metric for margin of confidence which selects the set of tweets that are most likely to be mis-classified. Adversarial debiasing involves training an adversary network to debias the classifier by including the adversary network’s loss. Similarly, (Mozafari et al., 2020) report results on hate speech detection from African-American and Standard American tweets. They re-weight instances based on the presence of phrases that they may highlight racial bias. They fine-tune BERT for the task. Finally, (Zhang et al., 2021) present an approach to reduce spurious correlation between two attributes: toxicity and African-American Vernacular English. They construct triplets of sentences where the first pair have the same toxicity label and the first and third have the same dialect label. The objective function of the model consists of a triplet loss over these triplets, and a disentanglement loss that ensures the masks for the true attributes are well-separated. Similarly, graphical models have been used to infer socio-cultural norms since they are closely associated with dialectic variations based on the language and cultural background of the speaker. (Moghimifar et al., 2023) present a Markov model to discover socio-cultural norms in emotion classification.
@@
Table 4. Sentiment Analysis Approaches for Dialectic Datasets.
@@
width = colspec = Q[290]Q[260]Q[580], hline1,7 = -0.08em, hline2 = -0.05em, Paper & Dialects Highlight
@@ 5.3. Morphosyntactic analysis
Morphosyntactic analysis deals with linguistic tasks such as POS tagging and morphological analysis. Morphological analysis has been found to be useful for sense disambiguation, particularly in low-resource settings (Khalifa et al., 2020). We now describe past work that deals with dialectic variations, as summarised in Table 5.
@@ 5.3.1. Classical approaches
(Habash and Rambow, 2006) is a seminal morphological analyser for dialects of Arabic called MAGEAD. Using morphological rewrite rules, they show how a morphological analyser can be adapted for dialects of a language. (Jørgensen et al., 2015) evaluate on a dataset of African-American Vernacular English and show that the then-prevalent POS taggers perform significantly worse. (Darwish et al., 2018) present a CRF-based POS tagger for dialects of Arabic. The POS tagger is trained on a small set of tweets using features derived from the dialects of interest. These features are progressive and negation particles.  (Eskander et al., 2016) adapt existing morphological analyzers to unseen dialects of Arabic by simulating the low-resource dialects.
@@ 5.3.2. Deep learning-based approaches
(Inoue et al., 2022) use CamelBERT trained on Modern Standard Arabic fine-tuned on dialect-specific datasets for morphosyntactic analysis. They observe that training using high-resource dialects helps low-resource dialects as well. In the context of Indic languages, (Bafna et al., 2023) explore POS tagging for 5 Indic dialects by focusing on Hindi-aware LLM adaptation via small dialectic monolingual corpora.
@@ 5.4. Parsing
Parsing involves the creation of semantic parse trees from text. Past work in parsing texts written in dialects of a language lies in three categories. The first category uses an existing parser on a dataset in a dialect of interest. The focus of such work is to create a baseline performance of popular parsers. The second category provides approaches to mitigate the bias of existing parsers towards texts in the dialect of a language. The third category creates a new parser for the dialect.

### v4

@@ 5.1.1. Shared tasks
Shared tasks have accelerated past work in dialect identification. These have been primarily led by Workshop on NLP for Similar Languages, Varieties and Dialects, also known as VarDial. The shared tasks are listed in Table 3.  Goutte et al. (2016) summarise the findings from the past versions of the shared task from 2014-2016. Similarly, the Nuanced Arabic Dialect identification (NADI) shared task is held annually. In the 2024 edition of NADI, a new task has been introduced to estimate the Arabic level of dialectness within Arabic sentences.
@@ 5.1.2. Datasets
Aji et al. (2022) report a dataset for several languages and dialects spoken in Indonesia. They observe that language identification works well for certain dialects (Ngoko-Central dialect of Javanese, for example). The paper also discusses code-mixing and orthography variations in these languages. Dunn (2019) reports dialect identification on 14 national varieties of English. He shows that cross-domain classification (CommonCrawl versus Twitter) also performs poorly. Cotterell and Callison-Burch (2014) present a multi-dialect, multi-genre corpus of news comments and tweets written in dialects of Arabic. The tweets are manually annotated for dialect identification on MTurk. Ramponi and Casula (2023a) present a benchmark dataset for dialects of Italian. The benchmark is named as DIATOPIT. There has been recent work on creating a corpus of Norwegian dialect (Barnes et al., 2021). Also, Alshutayri and Atwell (2018) present a large (200K+ instances) corpus for Arabic dialects and Standard Arabic. The data is sourced largely from tweets but also includes comments from newspapers, and Facebook. The data is also being annotated for dialect identification and contains 24K annotated documents. Recently, Talafha et al. (2024) introduced CASABLANCA, a large scale community-driven effort to collect and transcribe a multi-dialectal Arabic dataset, covering eight dialects for Arabic releasing $48$ hours of manually transcribed speech data including annotations for transcription, gender, dialect, and code-switching. Le and Luu (2023) present a parallel corpus for dialects of Vietnamese. Further, Dinh et al. (2024) propose a dialect identification, and speech recognition dataset, and fine-tuned models for for $63$ provincial dialects of Vietnamese with $102.5$ hours of audio, and $19000$ spoken utterances.
@@ 5.1.3. Feature-based approaches
We now highlight features used for dialect identification.
@@
    Phonological features: Phonological features are based on markers in the written scripts. Darwish et al. (2014) use lexical along with a lexicon of dialectal Egyptian words, morphological and phonological features in a random forest classifier to detect dialects of Arabic spoken in a geographical region.
@@
    Linguistic featuresDoğruöz and Nakov (2014) present a method to predict dialects of Turkish by using light verb constructions. They use a statistical classifier based on verb-based features (base word, verb order, affixes, etc.) for the task. Xie et al. (2024) discuss an approach to extract distinguishing lexical features of dialects by utilising interpretable dialect classifiers. With focus on varieties of Mandarin, Italian, and Low Saxon, this approach shows promising results on all varieties.
The combinations of the above set of features have also been reported. While Hanani and Naser (2020) work on the detection of dialects from speech, they also use word-level n-gram features. Salameh et al. (2018) perform fine-grained dialect identification for 25 dialects of Arabic, using Naïve Bayes classifier and word and character n-grams as features.
While dialect detection of Arabic has been explored in detail, the pre-deep learning work in the context of dialects of English is comparatively limited, although English is the predominant language for NLP researchLui and Cook (2013) is an early work in the detection of dialects of English. Specifically, the paper focuses on Australian, British and Canadian English. Their baseline is the LangID classifier (Lui and Baldwin, 2012) where dialects are treated as individual languages. They experiment with classifiers using features such as n-grams and POS-n-grams. This includes a distribution over function words and those in a vocabulary, akin to a clustering algorithm. Simaki et al. (2017) use linguistic, POS-tag-based and lexicon-based features.
@@ 5.1.4. Deep learning-based Approaches
Deep learning-based approaches for dialect classification span three alternatives: train embeddings to reflect dialectal variations, use end-to-end LLMs, or predict dialect as a result of inference over dialect features.
Embeddings in focusAbdul-Mageed et al. (2018) label tweets with 10 dialects of Arabic. The city is considered the dialectal granularity. The analysis compares dialectal variants by looking at word embeddings of words across different dialects. They use word2vec representations to show how dialectal words are captured. Goswami et al. (2020) build character-to-sentence embeddings to represent words of different dialects. Unsupervised loss is computed in order to generate clusters of representations. While they also test on language identification, the dialect identification part is done on Swiss German dialect. Jurgens et al. (2017) use a character-based seq2seq model to map dialects. The models used for language identification are RNNs with GRU. Criscuolo and Aluisio (2017) use character n-grams to identify language groups. This is followed by convolutional neural network-based dialect classifiers for each language group.
Fine-tuning LLMsRamponi and Casula (2023a) experiment with multiple models including statistical and neural. The fine-tuned AlBERTo model performs the best among umBERTo, mBERT and XLM-R. Obeid et al. (2020) present CAMeL: a python toolkit for Arabic language processing. It contains a dialect identifier that gives a distribution over multiple dialects. They use dialectal guidelines provided in  Elfardy and Diab (2012).
Detecting dialect features: d-eta-2021-learning introduce an approach for dialect classification using a novel multi-task approach that employs dialect feature detection. They train two multi-task learning-based approaches using a small number of minimal pairs. They evaluate the output based on 22 dialectal features based on Indian English and demonstrate that such models show the capability of learning to identify features with high accuracy. They show the efficacy of this task by applying it to dialect identification, and by providing a measure of dialect density.
@@ 5.2. Sentiment Analysis
Sentiment analysis is the NLU task of prediction of sentiment polarity of a text. Sentiment analysis encompasses several related tasks, such as sarcasm classification and target-specific sentiment analysis. We discuss past work in sentiment analysis along four directions: experiences from annotation (which highlights the challenge of dialects for sentiment analysis), dialect-aware models, dialect-invariant models, and, finally, de-biasing of sentiment analysis models as a post-processing step. Table 4 summarises approaches for sentiment analysis.
Early guessing for dialects: A recent advancement in dialect identification is early guessing (Kanjirangat et al., 2022). The approach detects a dialect for an incremental input.  Salloum and Habash (2022) also break the input down into its components. Specifically, they present an unsupervised approach that uses unsupervised dialect segmentation for machine translation.
@@ 5.2.1. Datasets & Annotation
Several datasets in Arabic sentiment analysis for dialects have been reported such as Moroccan (Oussous et al., 2020) and Levantine (Baly et al., 2019). Dialects can have an impact on annotation itself. Farha and Magdy (2022) show that dialect familiarity helps sarcasm annotation. Mdhaffar et al. (2017) create a dataset of 17000 Facebook comments labeled with sentiment in Tunisian dialect of Arabic. Assiri et al. (2018) present a sentiment-labeled lexicon of words in the Saudi dialect of Arabic, and use simple counting-based sentiment analysis. Husain et al. (2022) use weakly supervised labels for sentiment analysis of tweets in the Kuwaiti dialect of Arabic. The labels are then manually validated and updated.
@@ 5.2.2. Dialect-aware representations
Given the high degree of similarity between dialects, there is a high likelihood for models to make inferences in the same way for different dialects and thus explicitly modeling dialect awareness into models is important. However, this same degree of similarity makes this dialect aware modeling challenging. Farha and Magdy (2022) train BERT-based models for sarcasm detection on data annotated by either of the two groups: those familiar with the dialect and those not. They show that familiarity of dialect improves the quality of the models trained on such a dataset. As a result, representations that capture dialects have been used for sentiment analysis. Mdhaffar et al. (2017) present models based on SVM and multi-layer perceptron (MLP). Mulki et al. (2019) use a syntax-ignorant n-gram composition to create embeddings. The classifier model is a dense neural network that works on the addition of word embeddings, with a softmax at the end. Guellil et al. (2021) propose ‘one’ model for sentiment classification in different dialects of Arabic. They use transliteration to map dialects to Standard Arabic. The sentiment analysis model itself uses word2vec features with statistical classifier. Finally, Husain et al. (2022) present statistical models based on SVM along with Transformers-based models like BERT.
@@ 5.2.3. Incorporating dialect information in sentiment prediction
El Mekki et al. (2021) use domain adaptation for sentiment analysis of dialects. Using representations from a BERT encoder, they use two classifiers: sentiment classifier and dialect classifier. The output of the two is later combined for the overall prediction. While this is a two-channel approach, the representation used for the task has also been used to predict dialect of the language. One such example is Okpala et al. (2022) who present an approach for hate speech detection using African-American English. In order to do so, they re-train BERT with AAE tweets. Finally, adversarial training is needed to regulate the debiasing of the hate speech classifier. Specifically, the adversary takes the final representation learned by the hate-speech classifier, and learns to predict the dialect from it. Kaseb and Farouk (2022) present a dialect-aware approach for sarcasm detection called the SAIDS model. SAIDS uses MARBERT to detect dialect and sarcasm. Following that, MARBERT, along with sarcasm and dialect output, are used to detect sentiment. Evaluated on Arabic dialects, SAIDS uses backpropagation only for prediction with respect to the BERT base model. It does not flow through sentiment¡-¿sarcasm or sentiment¡-¿dialect.
@@ 5.2.4. De-biasing sentiment analysis models
Making sentiment analysis agnostic to dialects involves removing dialectal biases in the resultant models. A work of this nature is by Ball-Burack et al. (2021) who apply adversarial debiasing to resampled data for harmful tweet detection of tweets written in African-American English. Resampling of the data uses a metric for margin of confidence which selects the set of tweets that are most likely to be mis-classified. Adversarial debiasing involves training an adversary network to debias the classifier by including the adversary network’s loss. Similarly, Mozafari et al. (2020) report results on hate speech detection from African-American and Standard American tweets. They re-weight instances based on the presence of phrases that may highlight racial bias. They fine-tune BERT for the task. Finally, Zhang et al. (2021) present an approach to reduce spurious correlation between two attributes: toxicity and African-American Vernacular English. They construct triplets of sentences where the first two have the same toxicity label, and the first and the third have the same dialect label. The objective function of the model consists of a triplet loss over these triplets, and a disentanglement loss that ensures the masks for the true attributes are well-separated. Similarly, graphical models have been used to infer socio-cultural norms since they are closely associated with dialectal variations based on the language and cultural background of the speaker. Moghimifar et al. (2023) present a Markov model to discover socio-cultural norms in emotion classification. Harris et al. (2024) evaluate the zero-shot performance of speech recognition systems across different genders and across four US-based English dialects: SAE, AAVE, Chicano English, and Spanglish, release a labeled dataset of 13 hours of podcast audio, transcribed by speakers of the represented dialects.While past research has only dealt with African-American English, there may indeed be other dialects, which are considered aggressive and may result in sentiment analyzers producing biased output. This is significantly underexplored for dialects of other languages such as the Khariboli (Haryanvi group) dialect of Hindi (Yadav, 1974).
@@
Table 4. Sentiment Analysis Approaches for Dialectal Datasets.
@@
width = colspec = Q[290]Q[260]Q[580], hline1,8 = -0.08em, hline2 = -0.06em, Paper & Dialects Highlight
@@ 5.3. Morphosyntactic analysis
Morphosyntactic analysis deals with linguistic tasks such as POS tagging and morphological analysis, and has been found to be useful for sense disambiguation, particularly in low-resource settings (Khalifa et al., 2020). We now describe past work that deals with dialectal variations, as summarised in Table 5.
@@ 5.3.1. Classical approaches
Habash and Rambow (2006) is a seminal morphological analyser for dialects of Arabic called MAGEAD. Using morphological rewrite rules, they show how a morphological analyser can be adapted for dialects of a language. Jørgensen et al. (2015) evaluate on a dataset of African-American Vernacular English and show that the then-prevalent POS taggers perform significantly worse. Darwish et al. (2018) present a CRF-based POS tagger for dialects of Arabic. The POS tagger is trained on a small set of tweets using features derived from the dialects of interest. These features are progressive and negation particles.  Eskander et al. (2016) adapt existing morphological analyzers to unseen dialects of Arabic by simulating the low-resource dialects.
@@ 5.3.2. Deep learning-based approaches
Inoue et al. (2022) use CamelBERT trained on Modern Standard Arabic fine-tuned on dialect-specific datasets for morphosyntactic analysis. They observe that training using high-resource dialects helps low-resource dialects as well. In the context of Indic languages, Bafna et al. (2023) explore POS tagging for 5 Indic dialects by focusing on Hindi-aware LLM adaptation via small dialectal monolingual corpora.  Aepli and Sennrich (2022) propose improving cross-lingual transfer between closely related language varieties from the Finnic, West and North Germanic, and Western Romance language branches using character-level noise injection, and go on to show consistent improvements for POS tagging. Their approach is further applied to seven languages from three families and a total of eighteen dialects (Blaschke et al., 2023) with results showing improvements by varying the level of noise injected during the cross-lingual transfer.
@@ 5.4. Parsing
Parsing involves the creation of syntactic parse trees from text. Past work in parsing texts written in dialects of a language lies in three categories. The first category uses an existing parser on a dataset in a dialect of interest. The focus of such work is to create a baseline performance of popular parsers. The second category provides approaches to adapt existing parsers towards texts in the dialect of a language. The third category creates a new parser for the dialect.